Skip Navigation: Avoid going through Home page links and jump straight to content

MARS GLOBAL SURVEYOR STATUS REPORT

Wednesday, April 2 - DOY 092

Last Orbit Covered by this Report = 212

Total Aerobraking orbits accomplished = 180

Total Science Phasing orbits accomplished = 10

Apoapsis altitude = 17895 km
Apoapsis altitude decrease since start of aerobraking = 36131 km

Periapsis altitude = 170.7 km

Current Orbit Period = 11:39:43
Orbit Period decrease since start of aerobraking = 33:19:50

RECENT EVENTS:

The Science Phasing Orbit (SPO) operations are underway. The Mars Orbiter Laser Altimeter (MOLA), Mars Orbiter Camera (MOC), Magnetometer / Electron Reflectometer (MAG/ER), and Ka-Band Link Experiment (KaBLE) are all powered and returning data to Earth. In addition, the Radio Science (RS) team is collecting atmospheric science and gravity science data during the occultation period every orbit. The Thermal Emissions Spectrometer (TES) is currently powered on. The SPO profile for nominal orbits (not the Mars Pathfinder, Viking, or Cydonia imaging orbits ) is sequence driven and commands the S/C to nadir pointing (+Z axis to the center of Mars) 15 minutes prior to the closest point in the orbit (periapsis) and continues for 30 minutes. During this period the MOLA is within the altitude range for altimetry data collection and the MAG/ER collects data at a new sensitivity. Following the nadir period the S/C slews passing the MOC field of view across the planet. This roll-out primarily benefits the MOC imaging science collection and the rate and direction of the slew is set once per week. During the orbit all data is recorded onto the Solid State Recorders (SSRs) and is played back near the apoapsis period. One MOLA warming maneuver is executed during the orbit to maintain the instrument’s minimum temperature requirement.

All S/C subsystems report excellent health and performance. The attitude control performance remains un-eventful with perfect star processing performance. Reaction wheel momentum unloads is predicted to occur about once per 9 days during this period. The power subsystem continues its strong performance with no solar array current generation concerns. The batteries currently reach a maximum of 35.7% DoD and recharge in the expected time allotment. The telecommunications group reports continued excellent performance for both uplink and downlink components. Signal tracking through the occultation ingress and egress periods has been excellent. The S/C temperatures remain within limitations with the MOLA minimum temperature between 11°C and 12°C each orbit.

The regular weekly reset meeting was held today to review SPO operations and recommend new sequence parameters for the upcoming week. Parameters were presented and approved for special target imaging during the next week. Orbits 216, 217, 218 and 220 will employ special MOC roll-out modifications to attempt imaging of the Viking I lander area, the Viking II lander area, the Mars Pathfinder landing area and the Cydonia region. The first load of these orbit sequences will be prepared this afternoon and loaded tonight.

A command conference was held to approve the command files that will power the TES instrument and prepare it for data collection. These command will be uplinked tonight.

UPCOMING EVENTS:
(Note: MST = UTC-7 hours DOY090=3/31)
Solar Conjunction sequence start - April 30, 1998
Solar Conjunction - May 12, 1998

SPACECRAFT COMMANDING:

There were 9 command files radiated to the S/C during this period. The total files radiated since launch is now 1994. These commands were sent in support of the following activities:

MOLA mini-sequences and MOC NIPCs.

Nominal star catalog and ephemeris file update.